1. (a) From the definitions of probability density and flux, P(x,t) = 4*(x,t){(x,t) (:9 = 2 show that ƏP(x,t) at @j(x,t) Ox GCan anyone help me to solve this question on Quantum Mechanics about Schrodinger equation please?
Answer #1

9. Anscer 4 = 4 Exit P(x, f1 = 4*(x, f 14 (x, f) & j czt? = h ( 44 24 4 aut 7 911111111111 it dyrxt1 dt t² 22 2ucxth 3 4 ( xeajlat) Japcait) - ax a t o-o- o + V (27 ustup - vext y ý 1h Since nor is real, so so Vexi = vcxo* & 144*1 = 14441 aj (2447 +